A measured temperature on Fahrenheit scale is 200°F. What will this reading be on Celsius scale?

1
Expert's answer
2020-09-21T06:48:43-0400

Using the formula

"F= \frac{9}{5}C+ 32"

Substituting "F= 200" in the above equation, "200= \frac{9}{5}C+ 32"

"\implies 200-32=\frac{9}{5}C"

"\implies C= \frac{5}{9}\times 168=93.33"

So "200\u00b0F" is equivalent to "93.33\u00b0C"